# Squeezing information from radio surveys to probe the primordial Universe

### v1

@@ Abstract
A major goal of cosmology is to understand the nature of the field(s) which drove primordial Inflation. Through future observations, the statistics of large-scale structure will allow us to probe primordial non-Gaussianity of the curvature perturbation at the end of Inflation. We show how a new correlation statistic can significantly improve these constraints over conventional methods. Next-generation radio telescope arrays are under construction which will map the density field of neutral hydrogen to high redshifts. These telescopes can operate as an interferometer, able to probe small scales, or as a collection of single dishes, combining signals to map the large scales. We show how to fuse these these operating modes in order to measure the squeezed bispectrum with higher precision and greater economy. This leads to constraints on primordial non-Gaussianity that will improve on measurements by Planck, and out-perform other surveys such as Euclid. We forecast that $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})\sim 3$, achieved by using a small subset, $\mathcal{O}(10^{2}-10^{3})$, of the total number of accessible triangles. The proposed method can be used for any signal that peaks in squeezed configurations.
I Introduction
@@
Primordial non-Gaussianity (PNG) is one of the most promising observational avenues to probe the primordial Universe, due to its sensitivity to the physics of these early stages. Measuring PNG could reveal the field content and the amplitude of field interactions [1], as well as verify the quantum nature of the primordial fluctuations [2] that seed the large-scale structure (LSS) we observe today, as predicted by Inflation. This early Universe information can be extracted by characterizing how much the distribution of the primordial perturbations deviates from Gaussianity.
@@
We propose a novel method that utilises the functionalities of a radio telescope array in 21cm intensity mapping surveys, in order to measure the bispectrum of squeezed triangles, in an optimal and efficient way, while maximizing the PNG signal and avoiding regions contaminated by systematics. The squeezed bispectrum, with $k_{1}\approx k_{2}\gg\,k_{3}$, can be used to significantly improve current bounds on local PNG [3, 4, 5, 6]. Constraining $f_{\mathrm{NL}}^{\mathrm{loc}}$ provides a unique observational handle to rule out entire classes of inflationary models and enhance our knowledge of the early Universe [7].
@@
Currently, the tightest constraints on PNG are from cosmic microwave background (CMB) measurements [8]. The results imply that the primordial perturbation field is weakly non-Gaussian and hence most of the PNG information is contained in the primordial bispectrum, i.e. the lowest-order deviation from a Gaussian distribution. Most of the available information on local PNG from the CMB anisotropies bispectrum has been extracted. LSS surveys have the potential to significantly improve the current PNG bounds [9, 10], since a 3D dark matter tracer (e.g. galaxies) has considerably more bispectrum modes than the 2D CMB. However, most of these LSS modes are in the nonlinear regime, which is dominated by the gravitational evolution of cosmic structures. The subdominant PNG signal must be separated from the late-time non-Gaussian gravitational contributions with significant precision. Recent results from optical galaxy surveys [11, 12, 13, 14, 15], where an analytical perturbative model has been used (see e.g. [16]), are far from being competitive with the CMB bounds. The main reason is the complexity of accurate analytical models beyond the perturbative regime which is needed to extract the wealth of PNG information from the nonlinear scales. A promising alternative is to rely on numerical approaches, like simulation-based [17, 18, 19, 20] or field-level inference [21, 22].
@@
In this letter, we employ a perturbative analysis at high redshifts, which facilitates access to a large number of bispectrum modes in the perturbative regime. By also using high-density samples, we can significantly boost the PNG signal. To achieve this, we investigate upcoming high-redshift 21cm intensity mapping surveys [23, 24, 25, 26, 27, 28, 29, 30], focusing on the squeezed bispectrum, $k_{1}\approx k_{2}\gg\,k_{3}$. Due to its distinct triangle preference, which couples different scales, this shape is mostly orthogonal to the late-time gravitational part of the LSS bispectrum and therefore relatively easier to separate from the latter.
II PNG from 21cm intensity mapping
@@
In the post-reionization Universe most of the neutral hydrogen (HI) resides inside galaxies, thus detecting its 21cm emission line provides a tracer of the underlying matter distribution. The HI intensity mapping (IM) technique consists of making brightness temperature maps of the sky at different radio frequencies, by measuring the integrated emission line from many galaxies that reside within each pixel. The combined emission yields a large detectable signal, while probing very large volumes of the Universe from the epoch of re-ionisation up to present times. Such spectroscopic surveys are planned for example with the Square Kilometre Array Observatory (SKAO) [29] and the Hydrogen Intensity and Real-time Analysis eXperiment (HIRAX) [30] and will cover a wide sky area and deep redshift range ($0<z\lesssim 3$).
@@
An array of radio telescopes are set up to measure HI intensity in two distinct ways: in single-dish (SD) mode, by auto-correlating the signal of each dish individually or in interferometer mode (IF), where the signal from all dishes is cross-correlated. The first probes the large and linear scales of clustering, while the latter probes the intermediate to the very small and nonlinear scales (large $k$-mode values) with high angular resolution. Both modes of the instrument compliment each other in Fourier space, as shown in the schematic illustration of Fig. 1. Additionally the two functionalities differ in the behaviour of the instrumental noise, which is the dominant noise contribution [31]. Contrary to optical surveys, the shot noise can be negligible, up to small scales.

### v2

@@ Abstract
A major goal of cosmology is to understand the nature of the field(s) which drove primordial Inflation. Through future observations, the statistics of large-scale structure will allow us to probe primordial non-Gaussianity of the curvature perturbation at the end of Inflation. We show how a new correlation statistic can significantly improve these constraints over conventional methods. Next-generation radio telescope arrays are under construction which will map the density field of neutral hydrogen to high redshifts. These telescopes can operate as an interferometer, able to probe small scales, or as a collection of single dishes, combining signals to map the large scales. We show how to fuse these operating modes in order to measure the squeezed bispectrum with higher precision and greater economy. This leads to constraints on primordial non-Gaussianity that will improve on measurements by Planck, and out-perform other surveys such as Euclid. We forecast that $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})\sim 3$, achieved by using a small subset, $\mathcal{O}(10^{2}-10^{3})$, of the total number of accessible triangles. The proposed method identifies a low instrumental noise, systematic-free scale regime, enabling clean squeezed bispectrum measurements. This provides a pristine window into local primordial non-Gaussianity, allowing tight constraints not only on primordial non-Gaussianity, but on any observable that peaks in squeezed configurations.
1 Introduction
@@
Primordial non-Gaussianity (PNG) is one of the most promising observational avenues to probe the primordial Universe, due to its sensitivity to the physics of these early stages. Measuring PNG could reveal the field content and the amplitude of field interactions [1], as well as verify the quantum nature of the primordial fluctuations [2] that seed the large-scale structure (LSS) we observe today, predicted by Inflation. This early Universe information can be extracted by characterizing how much the distribution of the primordial perturbations deviates from Gaussianity.
@@
We propose a novel method that utilises the functionalities of a radio telescope array in 21cm intensity mapping surveys, in order to measure the bispectrum of squeezed triangles – in an optimal and efficient way, while maximizing the PNG signal and avoiding regions contaminated by systematics. The squeezed bispectrum, with $k_{1}\approx k_{2}\gg\,k_{3}$, can be used to significantly improve current bounds on local PNG [3, 4, 5, 6]. Constraining $f_{\mathrm{NL}}^{\mathrm{loc}}$ provides a unique observational handle to rule out entire classes of inflationary models and enhance our knowledge of the early Universe [7].
@@
Currently, the tightest constraints on PNG are from cosmic microwave background (CMB) measurements [8]. The results imply that the primordial perturbation field is weakly non-Gaussian and hence most of the PNG information is contained in the primordial bispectrum, i.e., the lowest-order deviation from a Gaussian distribution. Most of the available information on local PNG from the CMB bispectrum has been extracted. LSS surveys have the potential to significantly improve the current PNG bounds [9, 10], since a 3D dark matter tracer (e.g. galaxies) has considerably more bispectrum modes than the 2D CMB. However, most of these LSS modes are in the nonlinear regime, which is dominated by the gravitational evolution of cosmic structures. The subdominant PNG signal must be separated from the late-time non-Gaussian gravitational contributions with significant precision. Recent results from optical galaxy surveys [11, 12, 13, 14, 15, 16], where an analytical perturbative model has been used (see e.g. [17]), are far from being competitive with the CMB bounds. The main reason is the complexity of accurate analytical models beyond the perturbative regime which is needed to extract the wealth of PNG information from the nonlinear scales. A promising alternative is to rely on numerical approaches, like simulation-based [18, 19, 20, 21] or field-level inference [22, 23].
@@
In this work, we employ a perturbative analysis at high redshifts, which facilitates access to a large number of bispectrum modes in the perturbative regime. By also using high-density samples, we can significantly boost the PNG signal. To achieve this, we investigate upcoming high-redshift 21cm intensity mapping surveys [24, 25, 26, 27, 28, 29, 30, 31], focusing on the squeezed bispectrum. Due to its distinct triangle preference, which couples different scales, this shape is mostly orthogonal to the late-time gravitational part of the LSS bispectrum [32, 33, 9] and therefore relatively easier to separate from the latter.
2 PNG from 21cm intensity mapping
@@
In the post-reionization Universe most of the neutral hydrogen (HI) resides inside galaxies, thus detecting its 21cm emission line provides a tracer of the underlying matter distribution. The HI intensity mapping (IM) technique [34] consists of making brightness temperature maps of the sky at different radio frequencies, by measuring the integrated emission line from many galaxies that reside within each pixel. The combined emission yields a large detectable signal, while probing very large volumes of the Universe from the epoch of reionization up to present times. Such spectroscopic surveys are planned, for example with the Square Kilometre Array Observatory (SKAO) [30] and the Hydrogen Intensity and Real-time Analysis eXperiment (HIRAX) [31], which will cover a wide sky area and deep redshift range ($0<z\lesssim 3$).
@@
Radio telescope arrays can measure HI intensity in two distinct ways: in single-dish (SD) mode, by auto-correlating the signal of each dish individually and summing them, or in interferometer mode (IF), where the signals from all dishes are cross-correlated. The first probes the large and linear scales of clustering, while the latter probes the intermediate to the very small and nonlinear scales (large $k$-mode values) with high angular resolution. These modes of survey complement each other in Fourier space, see Fig.˜1. Additionally the two functionalities differ in the nature of the instrumental noise, which is the dominant noise contribution [35]. Contrary to optical surveys, the shot noise can be negligible, up to small scales.
In addition to the intrinsic instrumental scale limitations of an HI IM survey, there are known systematics. Astrophysical sources introduce nearly smooth foregrounds which overwhelm the very large-scale fluctuations along the line of sight ($k_{\parallel}\lesssim 0.01\,\text{$h$/Mpc}$) rendering them unusable [36, 34, 37, 38, 39, 40]. For an interferometer, an additional non-smooth foreground component, the foreground wedge, contaminates some of the modes transverse to the line of sight [37, 38, 41, 42, 43, 44, 43] – see Fig.˜1.
Figure 1: Schematic illustration of the ranges of radial and transverse wavenumbers that the two survey modes are sensitive to. Reproduced from [45]. © 2015. The American Astronomical Society. All rights reserved.
The presence of local PNG induces a large-scale scale-dependent behaviour in the power spectrum of a tracer [46, 47, 48, 49, 50, 51]. Hence, an LSS survey should be able to probe the largest scales of clustering in order to provide competitive power spectrum constraints on $f_{\mathrm{NL}}^{\mathrm{loc}}$. For the bispectrum, access to a wide scale range is essential, in order to probe enough squeezed triangles and boost the signal-to-noise ratio. Thus, a single-dish HI IM survey favours a power spectrum analysis [52, 53, 54, 55, 56], outperforming an interferometer survey. The reverse holds for the bispectrum $f_{\mathrm{NL}}^{\mathrm{loc}}$ constraints, where surveys designed for IF mode outshine those in SD mode [57].
3 The 21cm bimodal bispectrum
In this work, we propose a novel method to boost the capabilities of HI IM to tightly constrain the amplitude of local PNG, while excluding the contaminated $k$-modes from the analysis. The key point is that the two functionalities of the radio telescope probe complementary scales (see Fig.˜1). Combining the data from these two can provide access to a very wide scale range with high resolution: from the large and linear scales, as probed by the SD mode, to the very small and nonlinear ones, measured by the IF mode. This is optimal for a bispectrum analysis and especially for an observable that has a strong signal on squeezed triangles, like local PNG.
Squeezed configurations correlate different scales ($k_{1}\sim k_{2}\gg k_{3}$). Thus, cross correlating a density field $\delta^{\rm SD}(\bm{k})$, as probed by an SD mode survey, with two density fields $\delta^{\rm IF}(\bm{k})$ probed by a survey in IF mode, we can form an HI IM bispectrum that has access to a very large number of squeezed triangles. We define the bimodal bispectrum, synergising two HI IM surveys in two different operating modes, for an overlapping redshift range and sky area, as
$\langle\delta^{\rm IF}(\bm{k}_{1})\,\delta^{\rm IF}(\bm{k}_{2})\,\delta^{\rm SD}(\bm{k}_{3})\rangle=(2\,\pi)^{3}\,\delta_{\rm D}(\bm{k}_{123})\,B^{\rm SD\times IF}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})\;,$ (3.1)
where $\bm{k}_{123}\equiv\bm{k}_{1}+\bm{k}_{2}+\bm{k}_{3}$.
The fields $\delta^{\rm SD,IF}$ are fluctuations in the 21cm brightness temperature and their measurements in Fourier space depend on survey window functions $W^{\rm SD,IF}(\bm{k})$. Crucially, SD observations observe the actual sky, so that the fluctuation is a convolution in Fourier space, $\delta^{\rm SD}(\bm{k})=W^{\rm SD}(\bm{k})\star\delta(\bm{k})$}, whereas IF mode measures a Fourier transform of the sky, leading to a simple product, $\delta^{\rm IF}(\bm{k})=W^{\rm IF}(\bm{k})\,\delta(\bm{k})$}. (We omit the $z$-dependence for brevity.)
The window functions $W^{\rm SD,IF}(\bm{k})$ ensure that the measured $k$-modes are well within the intrinsic instrumental limits and away from regions affected by systematics (Fig.˜1), giving a contamination-free measurement of the synergy bispectrum. These limits include the fundamental limitations of each instrument [45], which depend on the size of the dishes, as well as the known astrophysical foregrounds and the wedge effect in the case of IF mode. The latter two exclude all scales that satisfy $k_{\parallel}<k_{\parallel{\rm min}}$ and $k_{\parallel}<A(z)\,k_{\perp}$ [57]. Foregrounds can be recovered with percent precision [61, 62, 63, 64, 65], thus we consider an optimistic cut-off value, $k_{\parallel{\rm min}}=0.005\,\text{$h$/Mpc}$.
The estimator for measuring the synergy bispectrum between the SD and IF modes is
$\hat{B}^{\rm SD\times IF}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})=\frac{1}{V_{s}\,V_{123}}\sum_{\bm{q}_{1}\in k_{1}}\sum_{\bm{q}_{2}\in k_{2}}\sum_{\bm{q}_{3}\in k_{3}}\delta_{\rm K}(\bm{q}_{123})\,\delta^{\rm IF}(\bm{q}_{1})\,\delta^{\rm IF}(\bm{q}_{2})\,\delta^{\rm SD}(\bm{q}_{3})\;,$ (3.2)
where $V_{s}$ is the volume of the survey within each redshift bin. The Kronecker delta ensures the formation of ‘fundamental triangles’ with sides $\bm{q}_{i}$, satisfying $\bm{q}_{123}=0$, that fall in the ‘triangular bin’ of width $\Delta k$ defined by the bin centres ($k_{1}$, $k_{2}$, $k_{3}$). The volume in Fourier space of the fundamental triangle bins is, in the thin shell limit $V_{123}=8\,\pi^{2}\,k_{1}\,k_{2}\,k_{3}\,(\Delta k)^{3}$ [66], where the bin size $\Delta k$ is taken to be the fundamental frequency of the survey, $\Delta k=k_{\rm f}=2\pi/V_{\rm s}^{1/3}$. For the extreme cases of flattened ($k_{1}=k_{2}+k_{3}$) and open triangles ($k_{1}\neq k_{2}+k_{3}$) this expression breaks down. Here we consider the analytic corrections for these configurations presented in [67].
The covariance of the bimodal bispectrum estimator is given by ${\sf C}_{ij}\equiv\big{\langle}\delta\hat{B}_{i}\,\delta\hat{B}_{j}\big{\rangle}={\sf C}_{ij}^{\rm G}+{\sf C}_{ij}^{\rm NG}$ , where $\delta\hat{B}_{i}=\hat{B}_{i}-\langle\hat{B}_{i}\rangle$, $i$ and $j$ indices denote the different triangle configurations, i.e. $B_{i}\equiv B(k_{1}^{i},k_{2}^{i},k_{3}^{i})$, and G, NG are the Gaussian and non-Gaussian contributions. In the thin shell limit ($k\gg\Delta k$), the Gaussian part can be written as
${\sf C}_{ij}^{\rm G}=\,\frac{(2\pi)^{6}}{V_{s}\,V_{123}}\,s_{123}\,\delta_{ij}\,P^{\rm IF}(\bm{k}_{1})\,P^{\rm IF}(\bm{k}_{2})\,P^{\rm SD}(\bm{k}_{3})\,$, (3.3)
where $s_{123}=6,2,1$ for equilateral, isosceles and scalene triangles respectively. The HI IM power spectrum $P^{\rm SD,IF}$ is the tree-level model, including the instrumental noise (see Appendix˜A for details).
The non-Gaussian part of the covariance is a complicated quantity to calculate analytically. However, in the case of squeezed configurations, like those considered in this work, it can be accurately approximated [^1] This analytic expression includes the bispectrum-bispectrum and power spectrum-trispectrum contributions. The latter is equivalent to the first in the case of squeezed configurations [67], thus the factor of 2 in front of Section 3. Moreover, the six-point function contribution is negligible for squeezed triangles and for the scales considered here, while the super-sample contribution is subdominant for all scales probed [68]. by [68, 67, 69]
${\sf C}_{ij}^{\rm NG}=2\,\frac{(2\,\pi)^{3}}{V_{s}\,V_{123}^{i}\,V_{123}^{j}}\Big{[}$ $\delta_{k_{1}^{i}k_{1}^{\,j}}\,U(k_{1}^{i},k_{1}^{\,j})\;B^{\rm SD\times IF}(\bm{k}_{1}^{\,j},\bm{k}_{2}^{i},\bm{k}_{3}^{i})\,B^{\rm SD\times IF}(\bm{k}_{1}^{i},\bm{k}_{2}^{\,j},\bm{k}_{3}^{\,j})$
$+8\,{\rm perm}\Big{]},$ (3.4)
where $U(k_{1}^{i},k_{1}^{j})=16\pi^{3}\,k_{2}^{i}\,k_{3}^{i}\,k_{2}^{\,j}\,k_{3}^{\,j}(\Delta k)^{5}$. For the extreme triangle configurations, we consider the analytic corrections presented in [67]. For $B^{\rm SD\times IF}$, the tree-level model is used and is presented in Appendix˜A. A well motivated assumption for the instrumental noise is that it is Gaussian [45]; in this case the non-Gaussian part of the covariance will not contain instrumental noise contributions.
4 Results
We present Fisher forecasts for the HI IM bimodal bispectrum on constraining the amplitude of local PNG. In a given redshift bin the Fisher matrix is
$F_{\alpha\beta}=\int{\rm d}^{3}\bm{k}_{i}\,\frac{\partial B^{\rm SD\times IF}(\bm{k}_{i})}{\partial\theta_{\alpha}}\,{\sf C}_{ij}^{-1}\,\frac{\partial B^{\rm SD\times IF}(\bm{k}_{j})}{\partial\theta_{\beta}}\,.$ (4.1)
Here $\theta_{\alpha}$ are the parameters of interest, $\bm{k}_{i}$ is an abbreviation for the three sides of the $i$th triangle, and $\int{\rm d}^{3}\bm{k}_{i}\equiv 1/(4\pi)\int_{-1}^{1}d\mu_{1}\int_{0}^{2\pi}d\phi\sum_{T}$, where $\mu_{1}=\hat{\bm{k}}_{1}\cdot\hat{\bm{z}}$ and $\phi$ characterize polar and the azimuthal orientations of a triangle relative to the line-of-sight direction, $\hat{\bm{z}}$. The forecasts are from the angle-averaged bispectrum. The sum $\sum_{T}$ is over all the Fourier space triangles that satisfy $k_{\rm min}\leq k_{3}\leq k_{2}\leq k_{1}\leq k_{\rm max}$.
We only consider a squeezed subset of these configurations, i.e. those that satisfy $k_{1}\approx k_{2}\geq 3\,k_{\rm 3}$, since they contain most of the signal for the local PNG. Although the bimodal bispectrum can benefit from all triangle shapes. The minimum value $k_{\rm min}=k_{\rm f}$, is the largest scale probed in the redshift bin by the survey, and the maximum value $k_{\rm max}$ corresponds to the smallest scale where the theoretical model is reliable. We use a conservative choice $k_{\rm max}(z)=0.75\,k_{\rm NL}(z)$ and an optimistic one $k_{\rm max}(z)=k_{\rm NL}(z)$, where $k_{\text{NL}}$ is given by the inverse square root of the one-dimensional velocity dispersion [60]. The analysis is thus confined within the perturbative regime, where the tree-level model offers good agreement with the numerical results [70, 71, 72, 73, 74]. The tree-level description is used to model the expected value of the bimodal estimator Eq.˜3.2, including the observational windows for the two operating modes. Any $\bm{k}$-mode outside of the window will result in the exclusion of the triangle from the Fisher matrix calculation.
Figure 2: Marginalized forecasts of $\sigma(f_{\mathrm{NL}}^{\mathrm{loc}})$ from the bimodal bispectrum (SKA $\times$ HIRAX) with and without foreground cut, for conservative and optimistic $k_{\rm max}$. Forecasts from each survey and from the summed signal without cross correlation (SKA + HIRAX) are also shown. Purple line is the Planck 2018 constraint [8].